OpenStudy (anonymous):

eliminate the parameter x=t^2+2, y=t^2-4 please explain the steps

OpenStudy (anonymous):

you can start by isolating the t^2, and plug in the other equation

OpenStudy (anonymous):

so... step 1: isolating the t^2 from either equation. I picked the second. you should get t^2 = y+4

OpenStudy (anonymous):

step 2: plug that in the first x = t^2 + 2 x = (y + 4) + 2

OpenStudy (anonymous):

there's your answer, and you can try to simplify to make it nice :)

hero (hero):

I was thinking to just simply write the expression for x - y x - y = (t^2 + 2) - (t^2 - 4) = t^2 - t^2 + 6 = 6 So x - y = 6

OpenStudy (anonymous):

@Hero That works too :). It just depends on the problem I guess

OpenStudy (anonymous):

There are answer choices to this practice problem, and @alias I used the problem x=t^2+2 and I got x-2=t^2 thus, y=x-2-4 and y=x-6 and that is one of the answer choices which is y=x-6, x> or equal to 1. Is this right? what does the greater than or equal to sign have to do with this problem?

OpenStudy (anonymous):

y=x-6 is right. And what's going on with the x> or equal to 1? Can you elaborate?

OpenStudy (anonymous):

Ohhh wait nevermind I answered my own question haahahah thank you! Can you help me with another problem? @alias

OpenStudy (anonymous):

lol, sure :)

OpenStudy (anonymous):

x= sqrt t , y=2t+5 How do I deal with the sqrt? @alias

OpenStudy (anonymous):

solve for t, then plug in the other one

OpenStudy (anonymous):

same procedure

hero (hero):

What happens if you take \(x = \sqrt{t}\) and then square both sides?

OpenStudy (anonymous):

the sqrt goes away? @Hero

hero (hero):

Yes and then you'd have \(x^2 = t\)

hero (hero):

From there, you could easily substitute for t in the second equation.

OpenStudy (anonymous):

Would the answer be y=2x^2+5? @Hero

hero (hero):

That's how I was imagining it.

OpenStudy (anonymous):

Yes! Thank you :D @Hero